A 120-V hair dryer has two settings: 850 W and 1250 W.

A) at which setting do you expect the resistance to be higher?After making a guess,determine the resistance at

B)the lower setting;and

C) the higher setting

Please show work.
Answer #1
A) P=V^2 /R, so assuming that the voltage is the same for bothhairdryers, the one with higher power should have a lowerresistance (they are inversely proportional)

B) P=V^2 /R   --> R= 120^2 /850W = 16.94 Ω

C) P=V^2 /R   --> R= 120^2 / 1250W = 11.52 Ω
Answer #2
A 120-V hair dryer has two settings: 850 W and 1250 W.

A) at which setting do you expect the resistance to be higher?After making a guess,determine the resistance at

B)the lower setting;and

C) the higher setting

Please show work.
power=voltage*current
but voltage =current*resistance
so current=V/R
hence power=(voltage)2/resistance
which clearly indicates
that power is inversely proportional to resistance
hence resistance is higher for 850W
B)
R=V*V/P=120*120/850=16.9411ohm
C)
R=V*V/P=120*120/1250=11.52ohm
Answer #3

The hair dryer has 2 settings R 850 W R 1250 W The relation between the voltage and the power, resistance is given as: T2The power is inversely related to the resistance. The greater the resistance, the lesser is the power Hence, the setting with